# Collisional whistler instability and electron temperature staircase in inhomogeneous plasma

### v1

@@ 8 Conclusion
Finally, we discuss the back-reaction of the collisional whistler instability on the plasma temperature profile (Sec. 7). The instability is able to modify the temperature profile via frictional heating and Ettingshausen heat flux so that total energy is conserved. Interestingly, there exists a regime in which the instability cools the plasma via friction rather than heats it; this regime necessarily occurs in the initial stages of the instability. The Ettingshausen heat flux is also capable of canceling a portion of the conductive heat flux when the intensity gradient of the collisional whistler instability is anti-aligned with the temperature gradient. In principle, the collisional whistler instability might be capable of strongly reducing the heat flux through these mechanisms, but for high-beta plasmas, a strong reduction is unlikely to occur in the manner envisioned here as this would require the wave-intensity profile to be essentially delta-shaped. Non-geometrical-optics behaviour, nonlinear effects, or even synergistic interplay with kinetic microinstabilities (as these quickly modify fluid transport coefficients away from the standard Braginskii expressions used here) might relax this conclusion, but dedicated simulations are required to investigate this further.
@@ Acknowledgements
The authors would like to thank Elijah Kolmes for insightful discussions. The work of NAL and, in part, of AAS was supported by STFC (grant number ST/W000903/1). The work of AAS was also supported in part by EPSRC (grant number EP/R034737/1) and the Simons Foundation via the Simons Investigator award. AFAB was supported by UKRI (grant number MR/W006723/1).
@@ A.1 No density fluctuations
The density equation (1a) demands that $n$ remain constant in time. Hence, there can be no fluctuating component to the density since that would require a nonzero time derivative. This is a simple consequence of quasineutrality: the ions are forced to remain static, so the electrons must remain static as well.
@@ A.2 No temperature fluctuations
00 0 $=\frac{c}{4\pi ne}(\nabla\times\widetilde{{\boldsymbol{\rm B}}})\cdot\left(\frac{3}{2}n\nabla T-T\nabla n+\overline{{\boldsymbol{\rm R}}}\right)+\frac{c}{4\pi ne}(\nabla\times\overline{{\boldsymbol{\rm B}}})\cdot\widetilde{{\boldsymbol{\rm R}}}-\nabla\cdot\widetilde{{\boldsymbol{\rm q}}}.$ (105)
@@
×$\nabla\times\overline{{\boldsymbol{\rm B}}}={\boldsymbol{\rm 0}},$ (106)
@@
(×$(\nabla\times\widetilde{{\boldsymbol{\rm B}}})\cdot\left(\frac{3}{2}n\nabla T-T\nabla n+\overline{{\boldsymbol{\rm R}}}\right)=0.$ (108)
@@ A.3 Verification of conditions for slab model
×$\nabla\times\overline{{\boldsymbol{\rm B}}}=0,\quad\nabla\times\widetilde{{\boldsymbol{\rm B}}}=\begin{pmatrix}-\widetilde{B}_{y}^{\prime}\\ \widetilde{B}_{x}^{\prime}\\ 0\end{pmatrix}.$ (110)
@@
where we have truncated at quadratic order in the fluctuation amplitude. Using the fact that $\nabla T$ is parallel to $\smash{\check{z}}$ (and thereby parallel to $\overline{{\boldsymbol{\rm B}}}$ and orthogonal to $\widetilde{{\boldsymbol{\rm B}}}$), that ×$\nabla\times\widetilde{{\boldsymbol{\rm B}}}$ is perpendicular to $\overline{{\boldsymbol{\rm B}}}$, and that
@@
(×$(\nabla\times\widetilde{{\boldsymbol{\rm B}}})\cdot\left(\frac{3}{2}n\nabla T-T\nabla n+\overline{{\boldsymbol{\rm R}}}\right)\propto(\nabla\times\widetilde{{\boldsymbol{\rm B}}})\cdot\smash{\check{z}}=0.$ (114)
@@ Appendix C Overview of Wigner–Weyl transform
Appendix D Wigner function bandwidth and quasi-eikonal fields
@@
$[\mathcal{I}^{\prime}(z)]^{2}=\mathcal{I}(z)\mathcal{I}^{\prime\prime}(z),$ (131)
@@
Appendix E Sufficient condition for positive temperature profile
@@
Appendix F Calculations pertaining to magnetization staircases
@@ F.4 Derivation of (72)
Appendix G Energy-conservation relations

### v2

@@ 9 Conclusion
Finally, we discuss the back-reaction of the collisional whistler instability on the plasma temperature profile (Sec. 8). The instability is able to modify the temperature profile via frictional heating and Ettingshausen heat flux so that total energy is conserved. Interestingly, there exists a regime in which the instability cools the plasma via friction rather than heats it; this regime necessarily occurs in the initial stages of the instability. The Ettingshausen heat flux is also capable of canceling a portion of the conductive heat flux when the intensity gradient of the collisional whistler instability is anti-aligned with the temperature gradient. In principle, the collisional whistler instability might be capable of strongly reducing the heat flux through these mechanisms, but for high-beta plasmas, a strong reduction is unlikely to occur in the manner envisioned here as this would require the wave-intensity profile to be essentially delta-shaped. Non-geometrical-optics behaviour, nonlinear effects, or even synergistic interplay with kinetic microinstabilities (as these quickly modify fluid transport coefficients away from the standard Braginskii expressions used here) might relax this conclusion, but dedicated simulations are required to investigate this further. In this sense, our work here should be considered an initial investigation in which a number of simplifications have been made to elucidate the basic physics at play and to obtain initial estimates of the various parameter dependencies. More detailed followup investigations can now be performed in which these simplifications are sequentially relaxed.
@@ Acknowledgements
The authors would like to thank Elijah Kolmes for insightful discussions.
Funding
The work of NAL and, in part, of AAS was supported by STFC (grant number ST/W000903/1). The work of AAS was also supported in part by EPSRC (grant number EP/R034737/1) and the Simons Foundation via the Simons Investigator award. AFAB was supported by UKRI (grant number MR/W006723/1).
Declaration of interests
The authors report no conflict of interest.
@@ A.1 No density fluctuations
The density equation (1a) demands that $n$ remain constant in time. Hence, there can be no fluctuating component to the density since that would require a nonzero time derivative.
@@ A.2 No temperature fluctuations
$0$ $=\frac{c}{4\pi ne}(\nabla\times\widetilde{{\boldsymbol{\rm B}}})\cdot\left(\frac{3}{2}n\nabla T-T\nabla n+\overline{{\boldsymbol{\rm R}}}\right)+\frac{c}{4\pi ne}(\nabla\times\overline{{\boldsymbol{\rm B}}})\cdot\widetilde{{\boldsymbol{\rm R}}}-\nabla\cdot\widetilde{{\boldsymbol{\rm q}}}$. (101)
@@
$\nabla\times\overline{{\boldsymbol{\rm B}}}={\boldsymbol{\rm 0}},$ (102)
@@
$(\nabla\times\widetilde{{\boldsymbol{\rm B}}})\cdot\left(\frac{3}{2}n\nabla T-T\nabla n+\overline{{\boldsymbol{\rm R}}}\right)=0.$ (104)
@@ A.3 Verification of conditions for slab model
$\nabla\times\overline{{\boldsymbol{\rm B}}}=0,\quad\nabla\times\widetilde{{\boldsymbol{\rm B}}}=\begin{pmatrix}-\widetilde{B}_{y}^{\prime}\\ \widetilde{B}_{x}^{\prime}\\ 0\end{pmatrix}$. (106)
@@
where we have truncated at quadratic order in the fluctuation amplitude. Using the fact that $\nabla T$ is parallel to $\smash{\check{z}}$ (and thereby parallel to $\overline{{\boldsymbol{\rm B}}}$ and orthogonal to $\widetilde{{\boldsymbol{\rm B}}}$), that $\nabla\times\widetilde{{\boldsymbol{\rm B}}}$ is perpendicular to $\overline{{\boldsymbol{\rm B}}}$, and that
@@
$(\nabla\times\widetilde{{\boldsymbol{\rm B}}})\cdot\left(\frac{3}{2}n\nabla T-T\nabla n+\overline{{\boldsymbol{\rm R}}}\right)\propto(\nabla\times\widetilde{{\boldsymbol{\rm B}}})\cdot\smash{\check{z}}=0.$ (110)
@@ Appendix C Overview of Wigner–Weyl transform
Appendix D Derivation of collisional whistler dispersion relation and growth rate via polar decomposition
As an alternative means of deriving the Hamiltonian (35), let us represent $\psi$ by its polar decomposition [^7] Note that this is traditionally the first step in the short-wavelength WKB approximation, whereafter one would introduce the additional assumption that $\theta$ varies more rapidly than $\mathcal{I}$. We shall not impose this latter constraint here, and thereby maintain an exact treatment.
$\psi=\sqrt{\mathcal{I}}\,\exp(i\theta).$ (122)
Thus, we see that the Hermitian and anti-Hermitian parts of the Hamiltonian identified in (35) via WWT-based methods emerges from the traditional approach as well. In fact, the evolution equation for the polar amplitude (125) is identical to that for the wave intensity given by (36). This is because, as discussed further in Appendix E, the final set of gradient terms in (125) encode the ‘non-eikonal’ bandwidth $\left\langle k_{z}^{2}\right\rangle-\left\langle k_{z}\right\rangle^{2}$ that can be combined with the term $\mathcal{D}_{A}\left(\partial_{z}\theta,z\right)$ to yield the Wigner-averaged growth rate $\left\langle\mathcal{D}_{A}\right\rangle$. The non-eikonal bandwidth being automatically contained in the Wigner-based formalism, rather than being a separate term that must be included in the evolution equations, is a theoretical advantage of that approach.
Appendix E Wigner function bandwidth and quasi-eikonal fields
@@
$\mathcal{I}^{\prime}(z)]^{2}=\mathcal{I}(z)\mathcal{I}^{\prime\prime}(z),$ (131)
@@
Appendix F Sufficient condition for positive temperature profile
@@
Appendix G Calculations pertaining to magnetization staircases
@@ G.4 Derivation of (68)
Appendix H Energy-conservation relations